OpenStudy (anonymous):

How do you workout this formula: a+b = 2xz; z

OpenStudy (lalaly):

do u want to solve for z?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

divide by 2x

OpenStudy (lalaly):

2xz=a+b z=(a+b)/2x

OpenStudy (anonymous):

I meant show the steps very well

OpenStudy (anonymous):

show it how I started it

OpenStudy (lalaly):

u just divide both sides by 2x thats it

OpenStudy (anonymous):

so divide 2x by a + b

OpenStudy (anonymous):

put a line then a 2x under a+b

OpenStudy (anonymous):

\begin{align*} a+b &= 2xz\\ \frac{a+b}{2x} &=\frac{\cancel{2x}z}{\cancel{2x}}\\ \frac{a+b}{2x} &=z \end{align*}

OpenStudy (lalaly):

the answer looks so pretty lol

OpenStudy (anonymous):

Ikr thanks so much

OpenStudy (anonymous):

but if x = 0 you've got a problem, if a + b = 0 z can be what you want, else z doesn't exist